# Source Code Foundation Models are Transferable Binary Analysis Knowledge Bases

### v1

@@ B.3 Importance and Use Scenarios of Function Name Recovery
B.4 Formal Definition of Precision and Recall Used by the SymLM Metrics
Formally, the precision and recall are defined as follows:
$Precision=\frac{\big{|}\!\big{|}T_{g}\cap T{p}\big{|}\!\big{|}}{\big{|}\!\big{|}T{p}\big{|}\!\big{|}}\quad Recall=\frac{\big{|}\!\big{|}T_{g}\cap T{p}\big{|}\!\big{|}}{\big{|}\!\big{|}T{g}\big{|}\!\big{|}},$
where $T_{g}$ is the token set of the groundtruth name, and $T_{p}$ the token set of predicted name.
Appendix C Prompts Used
@@
Appendix D User Study for Binary Summarization
The study leverages the same prompts used in the GPT4Evaluator (details in Appendix B.2), asking users to evaluate a piece of summary in terms of context relevance and functionality. The results are shown in Table 5. We can see that both CHRF and GPT4Evaluator are consistent with human preference, with GPT4Evaluator the most consistent metric with regard to human scores. Therefore, we use CHRF and GPT4Evaluator to evaluate the quality of binary summarizations.
In a pioneer study involving 120 random samples, we find that the score distribution is highly imbalanced. Only less than 5 samples have a score of 1 or 2 for both questions. To make the study more effective, the authors manually analyze 300 cases and construct a 60-sample question set with more balanced scores (based on the scores of authors). For each question, we make sure at least 3 users give their scores, and use the median scores as the results. We randomly shuffle summaries obtained with ProRec, a retriever, and directly prompting to prevent undesirable bias.
Table 5: Spearman correlation between human preference and auto-metrics. Columns 2–3 and 4–5 are for the context relevance questions and functionality questions, respectively. For each question, we report both the correlation and the p-value. A higher ‘correlation’ value and a smaller ‘p-value’ indicate a statistically stronger correlation.
@@ Appendix D User Study for Binary Summarization
Appendix E Case Study
We show two specific cases to illustrate the performance and limitation of ProRec. In Figure 11, we show a function that initializes a kay for encryption. Without any context information, GPT-3.5 summarizes the function with generic descriptions (e.g., “maniplate and transform the input data”). On the other hand, ProRec generates code snippets (only one of them is shown here) related to encryption keys. Provided with these code snippets, GPT-3.5 correctly summarizes the function as “perform cryptography operations”, and mentions several operations related to “key”. Although the summarization does not perfectly reflect the “initialization” purpose of this function, the description is clearer and more relevant to the context (i.e., key operations).
Figure 11: ProRec helps generate a summary that is more relevant to the source code context.
@@ Appendix E Case Study
Figure 12: ProRec helps generate a summary that is more relevant to the source code context. However, the recovered function name is different from source code.

### v2

@@ B.3 Importance and Use Scenarios of Function Name Recovery
B.4 Formal Definition of Precision, Recall, and F1 Used by the SymLM Metrics
Formally, the token-level precision ($P$), recall ($R$), and F1 are defined as follows:
$P(i)=\frac{\big{|}T_{g}^{(i)}\cap T_{p}^{(i)}\big{|}}{\big{|}T_{p}^{(i)}\big{|}}\quad R(i)=\frac{\big{|}T_{g}^{(i)}\cap T_{p}^{(i)}\big{|}}{\big{|}T_{g}^{(i)}\big{|}}\quad F1(i)=\frac{2\times P(i)\times R(i)}{P(i)+R(i)},$
where $T_{g}^{(i)}$ is the token set of the ground truth name for the $i$-th test case, and $T_{p}^{(i)}$ the token set of the $i$-th predicted name.
The precision, recall, and F1 scores for the entire test set are the average scores of individual scores across all test cases. Formally,
$P=\frac{1}{N}\sum_{i=1}^{N}P(i)\quad R=\frac{1}{N}\sum_{i=1}^{N}R(i)\quad F1=\frac{1}{N}\sum_{i=1}^{N}F1(i),$
where $N$ is the number of test cases.
Appendix C More Experiment Results and Analysis
C.1 Comparison with Supervised Binary Summarization Methods
Different from existing work CP-BCS [75] that supervisedly train a model for binary function summarization, ProRec does not require any superivsed training data for summarization and directly rely on LLM’s summarization ability. More importantly, CP-BCS’s summarization target is the docstring/comment of a function parsed from source code, which is not identical as the summarization targets in our experiments which are LLM summarizations from source code. For a fair comparison, we prepend the comments summarized by CP-BCS to the decompiled code as additional context for LLMs (gpt3.5-turbo-1106) to revise it into their own summarization styles, so that the final candidate summaries can be properly compared with reference source code summaries. Here, “+CP-BCS comment” means we augment the decompiled code with the comment for LLM to summarize. If we only evaluate the comments generated by CP-BCS, the CHRF drops to 5.44.
Table 5: CP-BCS generated comment-augmented binary summarization results.
CHRF G4-C G4-F
decompiled-only 30.4 3.6 3.8
+CP-BCS comment 29.0 3.0 2.8
We can see in Table 5 that CP-BCS comments have negative impacts on binary summarization results (based on gpt3.5-turbo-1106) on our test set, potentially due to the distribution difference between training and test data. In fact, we cannot easily adapt CP-BCS to this distribution since the training requires comments within the source code which do not exist in many functions in our training data. It is possible to distill summarization from LLMs, but the cost is high given the large amount of data. On the contrary, for ProRec data is less of a problem since all the compilable projects can be used to produce binary-source pairs that can be used for alignment.
C.2 Comparison with Black-box LLM as Prober
Leveraging black-box LLMs as probers is challenging because they are not heavily pre-trained on binary code and have limited understanding of it. ProRec addresses this through alignment training.
Table 6: Binary function name recovery results with self-probing on 100 examples.
[…]
decompiled-only 16.47 19.40 16.79
+retrieval 17.75 22.05 18.72
+ProRec 20.38 26.72 21.84
+self-probing 16.02 20.52 17.01
To demonstrate this empirically, we conduct experiments on binary function name recovery. We first prompt a black-box LLM (gpt3.5-turbo-1106) to translate decompiled functions into readable ones, sampling multiple results as diverse probed contexts. Using the same prompt as ProRec and the same LLM, we perform function name recovery with additional context. We call this method “self-probing”.
Table 6 shows the performance of self-probing (based on gpt3.5-turbo-1106) compared to other methods on 100 randomly sampled test data. We can see that self-probing performs slightly better than direct-prompting but is not comparable to retrieval-augmented recovery or ProRec.
Appendix D Prompts Used
@@
Appendix E User Study for Binary Summarization
The user study involved 12 participants. The participants are PhDs / PhD students that either have some background in reverse engineering or are experienced in C/C++/Rust programming. We ensured each summary is scored by at least 3 users, and use the median scores as the results. The questions in our user study were sampled from summaries generated by all three techniques (i.e., ProRec, the retrieval-augmented baseline, and direct prompting baseline).
The study leverages the same prompts used in the GPT4Evaluator (details in Appendix B.2) as the questionnaires, asking users to evaluate a piece of summary in terms of context relevance and functionality. As we already show the results of human judgment of different approaches in the main text, here we only show the results of Spearman correlation between human judgment and automatic metrics in Table 7. We can see that both CHRF and GPT4Evaluator are consistent with human preference, with GPT4Evaluator the most consistent metric with regard to human scores. Therefore, we use CHRF and GPT4Evaluator to evaluate the quality of binary summarizations.
Table 7: Spearman correlation between human preference and auto-metrics. Columns 2–3 and 4–5 are for the context relevance questions and functionality questions, respectively. For each question, we report both the correlation and the p-value. A higher correlation value and a smaller p-value indicate a statistically stronger correlation. Bold and underlined text indicate the best and second-best values, respectively, in each column.
@@ Appendix E User Study for Binary Summarization
Appendix F Case Study
In Figure 11, we show a function that initializes a key for encryption. Without any context information, GPT-3.5 summarizes the function with generic descriptions (e.g., “maniplate and transform the input data”). On the other hand, ProRec generates code snippets (only one of them is shown here) related to encryption keys. Provided with these code snippets, GPT-3.5 correctly summarizes the function as “perform cryptography operations”, and mentions several operations related to “key”. Although the summarization does not perfectly reflect the “initialization” purpose of this function, the description is clearer and more relevant to the context (i.e., key operations). We can see that retrieval results helps LLM generate a more context-relevant yet not functionally correct summary. That is because the datastore contains code snippets that are very similar to the query function (e.g., the function sp_256_ecc_recode in Figure 11-(d) is a crypto-related function that performs bit-wise operations). However, the retrieved function does not explicitly mention anything like “key”, which prevents the LLM recoverer to further guess the purpose of the function.
Figure 12 shows a more extreme case when RAG is less helpful than ProRec with no relevant function similar to the query function in the datastore. The query function pops an element from a queue. The retriever retrieves two snippets of code that have similar syntactic features (e.g., null pointer checks at the beginning; pointer accesses in the loop condition). By contrast, ProRec recognizes local semantic information such as getting an element from a queue, and atomic memory operations. Therefore, the probed code snippets are more relevant to program contexts even if the entire query function is not in the datastore.
Figure 11: ProRec helps to generate a summary that is more relevant to the source code context.
@@ Appendix F Case Study
Figure 12: ProRec can be more helpful than RAG in HOBRE tasks when no relevant function can be retrieved from the datastore.
Figure 13: ProRec helps to generate a summary that is more relevant to the source code context. However, the recovered function name is different from source code.